# Can LLMs Find Fraudsters? Multi-level LLM Enhanced Graph Fraud Detection

### v1

@@ Abstract.
copyright: acmcopyrightjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; October 27-31, 2025; Dublin, Irelandprice: 15.00isbn: 978-1-4503-XXXX-X/18/06ccs: Computing methodologies Machine learningccs: Information systems Web applications; Social networks
@@ 1. Introduction
Modern multimedia ecosystems are inherently manifested as complex multimodal graphs, such as social media platforms (Wei and Tyson, 2024), e-commerce systems (Lin et al., 2024), and video knowledge graphs (Xu et al., 2024), where heterogeneous nodes (users, products, multimedia content) interact through both structural connections and rich multimodal signals. The emerging paradigm of multimodal graph neural networks has demonstrated unprecedented capabilities in modeling these intricate relationships, particularly when augmented with Large Language Models (LLMs) that provide cross-modal semantic grounding for visual (Cai et al., 2024), textual (Wang et al., 2024a), and behavioral data (Wang et al., 2024b). However, the proliferation of sophisticated fraud patterns-ranging from fake multimedia reviews to AI-generated deceptive content-poses critical challenges to multimodal system integrity. These malicious campaigns often exploit semantic differences between graph topologies and unstructured modalities to undermine the reliability of multimedia recommendation and content understanding systems. Therefore, the development of robust multimodal graph fraud detection mechanisms is essential to ensure the veracity of cross-modal correlations while maintaining the complementary advantages of structural and semantic representations.
@@ 3.1. Preliminaries
The workflow of MLED is summarized in Algorithm 1. The algorithm takes the graph data, node features, node types, and relations as input, and outputs the predicted labels. It first generates initial node embeddings through a feature mapping layer. Then, it employs the type-level enhancer to distinguish fraudsters from benign entities and the relation-level enhancer to highlight critical relations for fraudsters. Finally, it fuses multi-level representations for fraud detection.
@@ 3.3. Type-level Enhancer
where $emb$ is the text embedding model, $\mathbb{T}$ denotes the set of node types, and $h^{LLM}_{t_{n}}\in\mathbb{R}^{\lambda}$ indicates the latent representation vector of type $t_{n}$ enhanced by the external knowledge extracted through LLM.
@@ 3.4. Relation-level Enhancer
where $emb$ is the text embedding model, $R$ denotes the set of relations, and $h^{LLM}_{r_{n}}\in\mathbb{R}^{\gamma}$ indicates the $\gamma$-dimensional latent representation vector of relation $r_{n}$ enhanced by the external knowledge extracted through LLM.
@@ 3.5. Fraud Detection
where $W_{f}$ is the trainable weight matrix, and $b_{f}$ is the bias vector. $\psi$ denotes the activation function softmax. For all the fraud detection methods, we utilize the cross entropy loss to train the model.
@@ 3.6. Time Complexity Analysis
Algorithm 1 Training procedure of MLED Framework
1:Input: Multi-relation graph $\mathcal{G}$, node features $\mathcal{X}$, node types $\mathbb{T}$, relations $R$.
2:Output: Predicted labels $\hat{y}_{t}$.
3:for each node $v_{t}\in\mathcal{V}$ do
4:     $h_{v_{t}}=\sigma(W_{h}\mathbf{x}_{t}+b_{h})$;
5:end for
6:for each type $t_{n}\in\mathbb{T}$ do
7:     Generate prompt $P(t_{n})$ ;
8:     Compute type-level embedding:
9:     $h^{LLM}_{t_{n}}=MLP(emb(P(t_{n})))$;
10:     Calculate type weight $\beta_{t_{n}}=\phi(W_{t}\cdot(h_{v_{t}}+h^{LLM}_{t_{n}})+b)$;
11:end for
12:Aggregate type embeddings: $z_{v_{t}}=\frac{1}{\mathbb{|T|}}(\sum_{t_{n}\in\mathbb{T}}\beta_{t_{n}}\cdot h^{LLM}_{t_{n}})$;
13:for each relation $r_{n}\in R$ do
14:     Generate prompt $P(r_{n})$;
15:     Compute relation-level embedding:
16:     $h^{LLM}_{r_{n}}=MLP(emb(P(r_{n})))$;
17:     Initial attention coefficient $\alpha_{r_{n}}=\sigma(W_{r}\cdot(h_{v_{t}}+h^{LLM}_{r_{n}}))$;
18:end for
19:Calculate relation weight $\delta_{r_{n}}=\frac{exp(\alpha_{r_{n}})}{{\textstyle\sum_{r_{n}\in R}}exp(\alpha_{r_{n}})}$;
20:Aggregate relation embeddings: $m_{v_{t}}=\sum_{r_{n}\in R}f_{att}(h^{LLM}_{r_{n}},\delta_{r_{n}})$;
21:Fuse multiple levels of representations:
22:$F_{v_{t}}=h_{v_{t}}+w_{tp}\cdot z_{v_{t}}+w_{re}\cdot m_{v_{t}}$
23:Fraud detection: $\hat{y}_{t}=\psi(W_{f}^{T}F_{v_{t}}+b_{f})$;
24:return $\hat{y}_{t}$;
@@ 4. Experiments
Table 1. Statistics of the datasets.
Dataset #Nodes Fraud(%) Relation #Edges
Amazon 11,944 821 (6.87%) U-P-U 175,608
U-S-U 3,566,479
U-V-U 1,036,737
YelpChi 45,954 6,677 (14.53%) R-U-R 49,315
R-T-R 573616
R-S-R 3,402,743
T-Finance 39,357
1,804
(4.58%)
Transaction 21,222,543
T-Social 5,781,065
174,280
(3.01%)
User Interaction 73,105,508
@@ 4.1. Experimental Setup
Datasets. We use the Amazon (McAuley and Leskovec, 2013), YelpChi (Rayana and Akoglu, 2015), T-Finance (Tang et al., 2022), and T-Social (Tang et al., 2022) datasets to evaluate the performance of MLED. Table 1 summarizes the statistics of these datasets. In the Amazon dataset, users are represented as nodes and there are three distinct relations between the nodes. The fraudsters are users with less than 20% helpful votes. To differentiate fraudsters from normal users, we label users with fewer than 20% helpful votes as fraudsters. The YelpChi dataset treats reviews as nodes, classifying them into two categories: spam and legit. The connections between nodes are also of three types. T-Finance aims to detect fraud accounts in transaction networks. The graph connections represent pairs of accounts with recorded transactions. T-Social aims to identify fraud users in social networks. Two users are connected if they have been friends for over three months.
@@ 4.3. Ablation Analysis
To investigate how the two level enhancers improve the performance of the proposed framework, we conduct an ablation analysis on the following variants of MLED. MLED_tl and MLED_rl denote the models that contain only the type-level enhancer and the relation-level enhancer, respectively. Figure 4 illustrates the experimental results of ConsisGAD in applying MLED and its variants. The results represent that the multi-level structure of framework outperforms both one-level variants in all metrics. The three evaluation metrics decreased by 0.9%, 0.92% and 0.55% on the Amazon dataset with only type-level enhancer, which decreased by 1.27%, 1.41% and 1.05% with only relation-level enhancer. Similarly, on the YelpChi dataset, the three evaluation metrics decreased by 0.73%, 2.33% and 2.17% with only type-level enhancer, while those decreased by 1.48%, 5.22% and 3.71% with only relation-level enhancer. It is noticeable that the type-level and relation-level enhancers play critical roles in MLED and their combination greatly benefits the fraud detection. It also demonstrates the effectiveness of the two proposed LLM enhancers in MLED. Moreover, the representation obtained from the type-level enhancer makes a greater contribution to our model. The results indicate that continuously widening the gap between fraudsters and benign entities is the key to detecting fraudsters.
@@
Figure 4. Ablation study results on various MLED variants.
Figure 5. Efficiency analysis on the Amazon and YelpChi datasets.
@@ 4.5. Visualization Analysis
4.6. Parameter Analysis
The proposed MLED framework contains two main hyperparameters, i.e., $\lambda$ and $\gamma$, which indicate the size of type-level representations and relation-level representations, respectively. To evaluate the robustness of MLED under varying hyperparameter configurations, we conduct a parameter sensitivity analysis on the Amazon and YelpChi datasets. This experiment aims to analyze the impact of the two hyper-parameters on fraud detection performance. Specifically, we vary the values of $\lambda$ and $\gamma$ within the range of [2, $\ldots$,64] and Figure 7 presents the experimental results. Based on the observations in the figure, we make several conclusions. Firstly, MLED tends to exhibit suboptimal performance when $\lambda$ and $\gamma$ are set to low values, e.g., 2 and 4. This emphasizes the crucial role of both representations in the MLED framework and highlights their effectiveness. Secondly, we can observe that excessively high values of $\lambda$ and $\gamma$ can also adversely affect performance, because they may mask the importance of the original node representations. Finally, MLED demonstrates relatively stable AUCROC performance across a wide range of $\lambda$ and $\gamma$ values, which proves its robustness.

### v2

@@ Abstract.
Multimodal Fusion; Graph Neural Network; Fraud Detection; Large Language Model
journalyear: 2025copyright: acmlicensedconference: Proceedings of the 33rd ACM International Conference on Multimedia; October 27–31, 2025; Dublin, Ireland.booktitle: Proceedings of the 33rd ACM International Conference on Multimedia ($M$ ’25), October 27–31, 2025, Dublin, Irelandisbn: 979-8-4007-2035-2/2025/10doi: 10.1145/3746027.3755245ccs: Computing methodologies Machine learningccs: Information systems Social networksccs: Information systems Data mining
@@ 1. Introduction
Modern multimedia ecosystems are inherently manifested as complex multimodal graphs, such as social media platforms (Chen et al., 2025b), e-commerce systems (Lin et al., 2024), and video knowledge graphs (Xu et al., 2024), where heterogeneous nodes (users, products, multimedia content) interact through both structural connections and rich multimodal signals. The emerging paradigm of multimodal graph neural networks has demonstrated unprecedented capabilities in modeling these intricate relationships, particularly when augmented with Large Language Models (LLMs) that provide cross-modal semantic grounding for visual (Cai et al., 2024), textual (Chen et al., 2025a), and behavioral data (Wang et al., 2024). However, the proliferation of sophisticated fraud patterns-ranging from fake multimedia reviews to AI-generated deceptive content-poses critical challenges to multimodal system integrity. These malicious campaigns often exploit semantic differences between graph topologies and unstructured modalities to undermine the reliability of multimedia recommendation and content understanding systems. Therefore, the development of robust multimodal graph fraud detection mechanisms is essential to ensure the veracity of cross-modal correlations while maintaining the complementary advantages of structural and semantic representations.
@@ 3.3. Type-level Enhancer
where embemb is the text embedding model, $\mathbb{T}$ denotes the set of node types, and $h^{LLM}_{t_{n}}\in\mathbb{R}^{\lambda}$ indicates the latent representation vector of type $t_{n}$ enhanced by the external knowledge extracted through LLM.
@@ 3.4. Relation-level Enhancer
where embemb is the text embedding model, $R$ denotes the set of relations, and $h^{LLM}_{r_{n}}\in\mathbb{R}^{\gamma}$ indicates the $\gamma$-dimensional latent representation vector of relation $r_{n}$ enhanced by the external knowledge extracted through LLM.
@@ 3.5. Fraud Detection
where $W_{f}$ is the trainable weight matrix, and $b_{f}$ is the bias vector. $\psi$ denotes the activation function softmax. For all the graph fraud detection methods, we utilize the cross entropy loss to train the model:
(9) $\mathcal{L}=-\sum_{t\in\mathcal{V}}y_{t}log\hat{y}_{t}+(1-y_{t})log(1-\hat{y}_{t}).$
@@ 4.1. Experimental Setup
Datasets. We use the Amazon (McAuley and Leskovec, 2013), YelpChi (Rayana and Akoglu, 2015), T-Finance (Tang et al., 2022), and T-Social (Tang et al., 2022) datasets to evaluate the performance of MLED. All the datasets consist of users and their associated comments. In the Amazon dataset, users are represented as nodes and there are three distinct relations between the nodes. The fraudsters are users with less than 20% helpful votes. To differentiate fraudsters from normal users, we label users with fewer than 20% helpful votes as fraudsters. The YelpChi dataset treats reviews as nodes, classifying them into two categories: spam and legit. The connections between nodes are also of three types. T-Finance aims to detect fraud accounts in transaction networks. The graph connections represent pairs of accounts with recorded transactions. T-Social aims to identify fraud users in social networks. Two users are connected if they have been friends for over three months.
@@ 4.3. Ablation Analysis
To investigate how the two level enhancers improve the performance of the proposed framework, we conduct an ablation analysis on the following variants of MLED. MLED_tl and MLED_rl denote the models that contain only the type-level enhancer and the relation-level enhancer, respectively. Figure 4 illustrates the experimental results of ConsisGAD in applying MLED and its variants. The results represent that the multi-level structure of framework outperforms both one-level variants in all metrics. The three evaluation metrics decreased by 0.9%, 0.92% and 0.55% on the Amazon dataset with only type-level enhancer, which decreased by 1.27%, 1.41% and 1.05% with only relation-level enhancer. Similarly, on the YelpChi dataset, the three evaluation metrics decreased by 0.73%, 2.33% and 2.17% with only type-level enhancer, while those decreased by 1.48%, 5.22% and 3.71% with only relation-level enhancer. It is noticeable that the type-level and relation-level enhancers play critical roles in MLED and their combination greatly benefits the fraud detection.
Figure 4. Ablation study results on various MLED variants with the SOTA method.
It also demonstrates the effectiveness of the two proposed LLM enhancers in MLED. Moreover, the representation obtained from the type-level enhancer makes a greater contribution to our model. The results indicate that continuously widening the gap between fraudsters and benign entities is the key to detecting fraudsters.
Figure 5. Efficiency analysis on the Amazon and YelpChi datasets.
@@ 4.5. Visualization Analysis
4.6. Parameter Analysis
The proposed MLED framework contains two main hyperparameters, i.e., $\lambda$ and $\gamma$, which indicate the size of type-level representations and relation-level representations, respectively. To evaluate the robustness of MLED under varying hyperparameter configurations, we conduct a parameter sensitivity analysis on the Amazon and YelpChi datasets. This experiment aims to analyze the impact of the two hyper-parameters on fraud detection performance. Specifically, we vary the values of $\lambda$ and $\gamma$ within the range of [2, $\ldots$,64] and Figure 7 presents the experimental results. Based on the observations in the figure, we make several conclusions. Firstly, MLED tends to exhibit suboptimal performance when $\lambda$ and $\gamma$ are set to low values, e.g., 2 and 4. This emphasizes the crucial role of both representations in the MLED framework and highlights their effectiveness. Secondly, we can observe that excessively high values of $\lambda$ and $\gamma$ can also adversely affect performance, because they may mask the importance of the original node representations.
@@ 4.6. Parameter Analysis
Finally, MLED demonstrates relatively stable AUCROC performance across a wide range of $\lambda$ and $\gamma$ values, which proves its robustness.
@@ 5. Conclusions
Acknowledgements.
The work is supported by the National Natural Science Foundation of China (No. 62272487 and 62433016).